Status Report

SOHO Daily Meeting Minutes for Friday, 10 May 2002

By SpaceRef Editor
May 10, 2002
Filed under , ,

* RHESSI target for today is AR 9934; target for tomorrow
is AR 9934.

* JOP104 target for today continues in the north east. Later
today SUMER will move south to catch AR 9934 as it reaches
the west limb.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: NOM HGA, VIRGO, SWAN, CELIAS, New Tracking Star

Planned Activities: RSL, VIRGO

Upcoming Operations: None

Ground Anomalies:
129/1622 14.25 minutes TM and CMD loss. D27. SCP failure. TM
is recoverable.
129/1643 47 minutes degraded TM. D27. RFI. Nonrecoverable.
130/0255 1.3 minute TM dropout. D46. TGC channel failure. Recoverable.

05/09 DOY 129 – 05/10 DOY 130

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Today – Spectral atlas, map of two active regions,
polarity reversal studies, lubrication study 20-22 UT and JOP136
support. Saturday and Sunday CDS will do weekend studies and JOP104.

SUMER: Nominal. JOP104 continues.

UVCS: Nominal. Coronal hole study in the southwest and long synoptic.
Weekend: coronal hole study in the southwest and mini-synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. Half-resolution 195 CME watch and synoptic.

MDI: Nominal. Full disk magnetograms and dopplergrams for continuous

TRACE: Nominal. Observing RHESSI region AR9934 in 195A.

SpaceRef staff editor.